OpenStudy (rational):

\[\large [3;6,\overline{1,4}] = 3+\dfrac{1}{6+\dfrac{1}{1+\dfrac{1}{4 + \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} =?\]

Parth (parthkohli):

\[{{\dfrac{1}{4 + \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} := ~ ~ t \] \[{{{4 + \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} = ~ ~ \frac{1}{t}\] \[{{{ \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} =\frac{1}{t} - 4 = t\]And so,\[1 - 4t = t^2\Rightarrow t^2 + 4t-1 =0 \Rightarrow t = \sqrt{5} - 2 \]

Parth (parthkohli):

Half of our job is done.

Parth (parthkohli):

\[3 + \frac{1}{6+t} = 3 + \frac{1}{\sqrt{5} + 4 }\]

OpenStudy (rational):

Neat !

Parth (parthkohli):

Is there some sort of trick when it comes to continuous fractions like these? Because I'm pretty sure what I showed above is just a standard-old-boring kind of solution.

OpenStudy (rational):

why didn't u let \[1+{{\dfrac{1}{4 + \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} := ~ ~ t\]

OpenStudy (rational):

these continued fractions always converge to an irrational number

OpenStudy (rational):

yeah they always converge

OpenStudy (perl):

is this correct? $$ \large {{{ \dfrac{1}{\color{red}1+\dfrac{1}{4+\cdots}}}}} =\frac{1}{t} - 4 = t $$

Parth (parthkohli):

Well, I suppose we could've done that, and substituted\[\frac{1}{6 + \frac{1}{t}}\] in the end.

OpenStudy (perl):

why is that equal to t, when t had a 4

Parth (parthkohli):

Ah, I read the fraction wrong.

Parth (parthkohli):

Thanks for pointing out. Let's make the modifications.\[{{{ \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} =\frac{1}{t} - 4 \]\[\dfrac{1}{1 + t} = \frac{1}{t} - 4\]I hope I'm right this time.

Parth (parthkohli):

\[t = t+1 - 4t(t+1)\]\[4t^2 + 4t -1=0\]

Parth (parthkohli):

\[t = \frac{\sqrt 2 -1}{2}\]

OpenStudy (rational):

im getting the same..

OpenStudy (perl):

looks good now :)

OpenStudy (rational):

\[\large [3;6,\overline{1,4}] = 3+\dfrac{1}{6+\dfrac{1}{1+\dfrac{1}{4 + \dfrac{1}{1+\dfrac{1}{4+\cdots}}}}} = 3+\dfrac{1}{6+\dfrac{1}{1+t}}\]

OpenStudy (rational):

should end up with \[\dfrac{14-\sqrt{2}}{4}\]

OpenStudy (rational):

the reverse of this problem is more interesting : represent \(\sqrt{2}\) as infinite continued fraction

OpenStudy (rational):

there seems to be a nice method to write any irrational number as a continued fraction \[\sqrt{2} = [1;\overline{2}]=1+\dfrac{1}{2+\dfrac{1}{2+\cdots}}\]

OpenStudy (rational):

http://gyazo.com/43e10aaa20413f2ba38cd9f2e8b6b8a6

Parth (parthkohli):

The continued fraction of \(\sqrt[3]{2}\), on the other hand, doesn't really seem to help.

Parth (parthkohli):

It's like an "irrational" continued fraction expansion, y'know.

OpenStudy (rational):

i mean the terms in continued fraction form dont seem to be periodic

OpenStudy (rational):

http://gyazo.com/48db46211dcfe4fe5dc9d7115f773841

OpenStudy (rational):

wolfram seem to know how to work these continued fraction forms quickly !

OpenStudy (rational):

this doesnt look periodic too http://gyazo.com/0c07cf4c190f6f341b9fff30c33ea2b9

OpenStudy (rational):

see http://gyazo.com/512d974365626a6baa125e12ce413d5b

Parth (parthkohli):

Yeah, I can understand that since \(\pi \) isn't algebraic.

OpenStudy (rational):

\(e\) has some pattern to it though

Parth (parthkohli):

Surprising...

OpenStudy (rational):

\[e = [2;\overline{1,2n,1}]\] where \(n=1,2,3,\ldots\)

OpenStudy (rational):

its not exactly periodic but it has that surprising pattern !

OpenStudy (perl):

and phi has a very nice pattern

OpenStudy (perl):

very simple i mean

OpenStudy (perl):

what is the continued fraction for cube root of 2

OpenStudy (rational):

\(\sqrt[3]{2}\) doesn't look periodic http://www.wolframalpha.com/input/?i=continued+fraction+2%5E%281%2F3%29

OpenStudy (rational):

may be not all irrational numbers are periodic in contined fraction representation

OpenStudy (perl):

right, but at least the square root of integers are

OpenStudy (rational):

\[\phi = [1;\overline{1}] = 1+\dfrac{1}{1+\dfrac{1}{1+\cdots}}\] i think this is the most nice looking continued fraction

OpenStudy (perl):

yes its surprising, and e is nice as well

OpenStudy (rational):

interesting.. is it easy to prove square roots of integers are periodic in continued fraction representation

OpenStudy (rational):

all the examples i saw so far are periodic